OpenStudy (anonymous):

Solve the equation for x: -3x2 + x + 4 = 0 Answer a. -4/3, -1 b. 4/3, 1 c. 4/3, -1 d. -4/3, 1

OpenStudy (anonymous):

\[x= - 1, 4/3\]

OpenStudy (anonymous):

hmm it does not line up with your choices :\ what did i do wrong, let me check

OpenStudy (anonymous):

No its C i think

OpenStudy (anonymous):

Yea it looks like it, Ugh i cant find out what i messed up

OpenStudy (anonymous):

Maybe animalain can help :)

OpenStudy (anonymous):

This one will factor.\[-3x ^{2}+x+4=-3x ^{2}+4x-3x+4=x(-3x+4)+(-3x+4)=(x+1)(-3x+4)\]

OpenStudy (anonymous):

Roots follow from this, x=-1 or x=4/3

OpenStudy (anonymous):

gopeder, the answer is c, but you have the order reversed.

OpenStudy (anonymous):

I guess I do, too....